OpenStudy (anonymous):

Seven subtracted from 3 times a number is equal to 9 more than 5 times the number. What is the number?

OpenStudy (anonymous):

3x-7=5x+9 -2x-7=9 -2x=16 -8

OpenStudy (anonymous):

thank you :D